From GlueXWiki
Jump to: navigation, search

August 21, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status Curved Tracks (Mike, Beni)
  3. FDC status scattering event, horizontal shower event (Lubomir)
  4. Electronics (Chris, Nick)
  5. Engineering (Bill)
    • Chiller
    • Cathode scan (Luke)
  6. Commissioning Wiki Page, CDC, FDC commissioning plan (all)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Sean (NU), Beni, Mike, Luke, Eugene, Chris, Nick, Dave, Bill, Fernando, Sergey, Vlad and Lubomir (JLab).


- Beni: testing the latest fADC125 firmware released, works except something goes wrong in the FPGA when setting event number. ROCFDC3 has problems will be excluded for the next run. We want to take data with the magnet on. The top scintillator paddle was moved as high as possible to reduce the filed on the PMTs. Actually, after the meeting Beni and Elton were able to get signals from the scintillator (after increasing the HV) and started the DAQ (FDC+CDC) with the coincidence trigger.

- Discussions about horizontal trigger. We need to install it back. The upstream paddles can be installed when the target is out. Instead of the downstream paddles one can use the TOF, but it requires the GTP.

CDC update

- Mike showed curved tracks (linked) in the CDC at 600A magnetic field. Some local curvatures seen can be explained if the drift time is taken into account.

- Fernado: additional ground on the CDC aluminum ring to which the HV is connected has been installed. Expect to reduce the noise significantly. Grounding the HVBs directly will be done as a last resort and it will be done in a different way through the grounding lugs.

FDC update

- Lubomir: by applying timing cuts on the hits from the two cathodes and the wires in a cell, one can reduce the threshold down to 50 fADC units (~4mV). As a result we see now most of the "missing" cathode channels, that have problems with the conductive tape. Of course the resolution will be deteriorated but one can still use this information for the pattern recognition. Couple of interesting events are linked above.


- Bill: problems with the chiller on Monday, it was not regulating the temperature. Tom adjusted the thermostatic valve and it started regulating again. Bill discussed with the company the chiller problems. They say the thermostatic valve must be OK, but the new compressor installed is not quite the same as the original one. They don't have spare units and they don't make visits to fix the chiller. If the problems continue, either we have to swap the compressor, or send the chiller to Indiana for repair, or get a new one from Japan. The plan for now is to install the heat exchanger (maybe tomorrow) and continue monitoring.

- On Monday when the chiller failed, Wesley Moore was working on the epics communication with the device. According to Hovanes he changed the set temperature to a lower value for as short time, then the chiller failed. The plan is just to read the epics variables, and not to write anything to the device.

- Luke: we installed the encoder and started measurements at the last position, but in the middle the encoder stopped working. In addition motor got disconnected from the stage. After the meeting the motor was fixed but there were still problems with the encoder. Luke will continue with measurements at two positions. The data with the encoder is enough to compare the two measurement methods.


- Curtis presented a plan for the CDC commissioning linked above. When aligning the photon beam several inner HVBs will be supplied by Bertan modules so that one can see the currents which may help in the beam alignment. Then the plan is to have some initial data with magnet on for a short time, then for a longer time with the magnet off, and then with the magnet on again. They are concerned that the timing algorithm on the fADC125 may not be ready for the run. Eugene: while working on the photon beam the magnet will stay at 300A which is safe; this time can be used for the initial data collection.

- As Bill proposed on the last meeting it will be good to survey the FDC without magnetic field (planned already for September) and then to monitor one of the reflector while the magnet goes up. Eugene: the best would be to rump up the magnet when the Hall is opened in September and at the same time the survey can be done.


- Beni: Scot put alcohol in the bubblers, we will monitor the chamber behavior.